Circular No. 9097

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 S1 = P/2001 Q10 (GIBBS) R. Matson (Irvine, CA) has identified astrometric observations of this

comet (cf. IAUC 9074) on NEAT 1.2-m reflector frames taken on 2001 Aug. 23 and 24 at Haleakala and on 2001 Sept. 24 and Oct. 27 at Palomar.

2001 UT  α_{2000}  δ_{2000}  Mag.
Aug. 23.60821  2^{h}14^{m}09.^{s}59  − 6^{o}22′20′′.8  20.0
24.61243  2 14 47.80  − 6 35 40.8  20.2
Sept. 24.48475  2 22 42.84  −14 38 01.7  19.8
24.50631  2 22 42.65  −14 38 22.6  19.3
Oct. 27.27596  2 09 16.92  −22 00 18.3  19.3
27.28631  2 09 16.49  −22 00 23.6  19.9
27.29669  2 09 16.13  −22 00 28.7  19.7

The following revised orbital elements by B. G. Marsden link the 2001 and 2009 apparitions; 2001 elements and residuals appear on MPEC 2009-X01.

Epoch = 2009 July 28.0 TT

T = 2009 Aug. 4.1741 TT  ω = 224.^{o}0034
e = 0.378003  Ω = 157.9798 }2000.0
q = 2.440203 AU  i = 26.1102

a = 3.923174 AU n^{o} = 0.1268375 P = 7.77 years

COMET P/2009 W1 (HILL) Additional astrometry has shown that this comet (cf. IAUC 9095) is

of short period; orbital elements from MPEC 2009-W147:

T = 2009 Sept.26.669 TT  ω = 52.^{o}372
e = 0.33828  Ω = 56.066 }2000.0
q = 2.97739 AU  i = 14.625
a = 4.49949 AU n^{o} = 0.103266 P = 9.54 years

V496 SCUTI Visual magnitude estimates (cf. IAUC 9095): Nov. 19.726 UT, 7.2

(K. Hornoch, Vohancice, Czech Republic); 21.719, 7.4 (P. Schmeer, Bischmisheim, Germany); 22.067, 7.5 (C. Hergenrother, Tucson, AZ, U.S.A.); 25.672, 7.5 (K. Sárneczky, Budapest, Hungary); 26.708, 7.4 (W. Vollmann, Vienna, Austria); 27.065, 7.5 (Hergenrother); 28.683, 7.2 (Sárneczky).

2009 December 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green